# Exact Bayesian Inference for Markov Switching Diffusions

### v1

@@ Appendix C Approximate MCMC Algorithm
where the latter is due do weak convergence of the Milstein approximation. Conversely, at higher imputation rates, such an imputation scheme negatively affects mixing, as examined in detail in roberts2001inference. As in the exact algorithm, we address this by switching to the non-centered parameterization. Defining $\bar{z}_{(\dot{\tau},\ddot{\tau})}=\zeta_{\theta}(\bar{x}_{(\dot{\tau},\ddot{\tau})};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$,
$\begin{split}&\bar{\pi}(v_{\ddot{\tau}},\bar{z}_{(\dot{\tau},\ddot{\tau})}|v_{\dot{\tau}},y_{\dot{\tau}},\theta)\\ &=\lvert\eta_{\theta}^{\prime}(v_{\ddot{\tau}})\rvert\prod_{z_{t}\in\bar{z}_{(\dot{\tau},\ddot{\tau})}}\lvert\partial_{z_{t}}\zeta_{\theta}^{-1}(z_{t};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\rvert\\ &\quad\prod_{(\dot{u}\sim\ddot{u})\in u_{(\dot{\tau},\ddot{\tau})}}\operatorname{\mathrm{N}}\left[\zeta_{\theta}^{-1}(z_{\ddot{u}};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}});\begin{subarray}{c}\zeta_{\theta}^{-1}(z_{\dot{u}};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})+(\ddot{u}-\dot{u})\delta_{\theta}(\zeta_{\theta}^{-1}(z_{\dot{u}};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}}),y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}}),\\ (\ddot{u}-\dot{u})\rho_{\theta}^{2}(y_{\dot{\tau}})\end{subarray}\right],\end{split}$ (93)
@@ Appendix D Poisson Coin Algorithm
for various paths $f:[0,\omega]\to[{f}^{\downarrow},{f}^{\uparrow}]$. This is addressed by the Poisson coin algorithm of (beskos2005exact). Notice that $f$ has to be upper bounded at ${f}^{\uparrow}$ in order to implement the Poisson coin algorithm. The main insight is that if we can construct and assess a tractable event $E$ such that $\operatorname{\mathrm{Pr}}\left[E\right]=p$, evaluation of $p$ is not necessary to flipping $p$-coins. We recall that a homogeneous Poisson process $\Psi$ on $\mathbf{R}^{d}$ is defined as a point process which satisfies
@@
We note that the complementary event $\{|\operatorname{\mathrm{epi}}\left[f-{f}^{\downarrow}\right]\cap\Psi|>0\}$ can often be ascertained without simulating the entire Poisson process, since only one point must fall into the epigraph - indeed, we recommend simulating $\Psi$ in slices from the bottom up, e.g. $[0,\omega]\times[0,({f}^{\uparrow}-{f}^{\downarrow})/2]$ and $[0,\omega]\times[({f}^{\uparrow}-{f}^{\downarrow})/2,{f}^{\uparrow}-{f}^{\downarrow}]$, and checking for each slice whether one of the points falls into the epigraph. If so, the simulation of the remaining slices can be skipped, since the Poisson coin is already known to be 0.

### v2

@@ Appendix C Approximate MCMC Algorithm
where the latter is due do weak convergence of the Milstein approximation. Conversely, at higher imputation rates, such an imputation scheme negatively affects mixing, as examined in detail in [40]. As in the exact algorithm, we address this by switching to the non-centered parameterization. Defining $\bar{z}_{(\dot{\tau},\ddot{\tau})}=\zeta_{\theta}(\bar{x}_{(\dot{\tau},\ddot{\tau})};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$,
$\begin{split}&\bar{\pi}(v_{\ddot{\tau}},\bar{z}_{(\dot{\tau},\ddot{\tau})}|v_{\dot{\tau}},y_{\dot{\tau}},\theta)\\ &\quad=\lvert\eta_{\theta}^{\prime}(v_{\ddot{\tau}})\rvert\prod_{z_{t}\in\bar{z}_{(\dot{\tau},\ddot{\tau})}}\lvert\partial_{z_{t}}\zeta_{\theta}^{-1}(z_{t};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\rvert\\ &\qquad\times\prod_{(\dot{u}\sim\ddot{u})\in u_{(\dot{\tau},\ddot{\tau})}}\operatorname{\mathrm{N}}\left[\begin{subarray}{c}\zeta_{\theta}^{-1}(z_{\ddot{u}};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}});\\ \zeta_{\theta}^{-1}(z_{\dot{u}};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})+(\ddot{u}-\dot{u})\delta_{\theta}(\zeta_{\theta}^{-1}(z_{\dot{u}};y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}}),y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}}),\\ (\ddot{u}-\dot{u})\rho_{\theta}^{2}(y_{\dot{\tau}})\end{subarray}\right],\end{split}$ (91)
@@ Appendix D Poisson Coin Algorithm
for various paths $f:[0,\omega]\to[{f}^{\downarrow},{f}^{\uparrow}]$. This is addressed by the Poisson coin algorithm of [7]. Notice that $f$ has to be upper bounded at ${f}^{\uparrow}$ in order to implement the Poisson coin algorithm. The main insight is that if we can construct and assess a tractable event $E$ such that $\operatorname{\mathrm{Pr}}\left[E\right]=p$, evaluation of $p$ is not necessary to flipping $p$-coins. We recall that a homogeneous Poisson process $\Psi$ on $\mathbf{R}^{d}$ is defined as a point process which satisfies
@@
We note that the complementary event $\{|\operatorname{\mathrm{epi}}\left[f-{f}^{\downarrow}\right]\cap\Psi|>0\}$ can often be ascertained without simulating the entire Poisson process, since only one point must fall into the epigraph - indeed, we recommend simulating $\Psi$ in slices from the bottom up, e.g. $0,\omega]\times[0,({f}^{\uparrow}-{f}^{\downarrow})/2]$ and $0,\omega]\times[({f}^{\uparrow}-{f}^{\downarrow})/2,{f}^{\uparrow}-{f}^{\downarrow}]$, and checking for each slice whether one of the points falls into the epigraph. If so, the simulation of the remaining slices can be skipped, since the Poisson coin is already known to be 0.
Appendix E Pseudocode Specification
Below we provide a pseudocode specification that consolidates the description of the Gibbs sampler in Section 3 and its generalization in Appendix B. Since their exact implementation goes beyond the scope of this paper, we do not fully specify the primitives $\texttt{construct-hidden-coin}(\theta,y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$ and $\texttt{construct-param-coin}(\theta^{\dagger},\theta,y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$. The former returns a function that takes a representation of $z_{(\dot{\tau},\ddot{\tau})}$ and returns a coin toss of probability $q_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$ and the updated representation of $z_{(\dot{\tau},\ddot{\tau})}$, including any further information revealed retrospectively. The latter takes a representation of $z_{(\dot{\tau},\ddot{\tau})}$ and returns coin tosses of probability $\exp\{-\int_{\dot{\tau}}^{\ddot{\tau}}\xi_{\theta^{\dagger},\theta}(z_{t},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})\mathop{}\!\mathrm{d}t\}$ and the updated representation of $z_{(\dot{\tau},\ddot{\tau})}$. Their implementation with the Poisson coin algorithm is sketched out in Section 3.1 and Supplement D, and described in further detail e.g. in [22]. The notation below uses the product operator to aggregate multiple coin toss functions into a single coin toss function with probability equal to the product of the constituting probabilities, e.g. $\texttt{construct-hidden-coin}(\theta,y_{\tau_{0}},v_{\{\tau_{0},\tau_{1}\}})\times\texttt{construct-hidden-coin}(\theta,y_{\tau_{1}},v_{\{\tau_{1},\tau_{2}\}})$ results in a function that takes a representation of $z_{(\tau_{0},\tau_{1})}\cup z_{(\tau_{1},\tau_{2})}$ and returns a coin toss of probability $q_{\theta}(z_{(\tau_{0},\tau_{1})},y_{\tau_{0}},v_{\{\tau_{0},\tau_{1}\}})\times q_{\theta}(z_{(\tau_{1},\tau_{2})},y_{\tau_{1}},v_{\{\tau_{1},\tau_{2}\}})$ as well as the updated representation.
Algorithm 2 Update for $\pi(\theta,\lambda,v_{r},z,y|v_{s})$ (See Section 3).
function update-all($\theta,\lambda,v_{\tau},y,z$)
  $v_{r^{*}}^{*},z^{*},y^{*}\leftarrow\textsc{update-hidden}(\theta,\lambda,v_{r},v_{s},y,z)$
  $\theta^{*},z^{*}\leftarrow\textsc{update-param}(\theta,v_{\tau^{*}}^{*},z^{*},y^{*})$
  $\lambda^{*}\leftarrow\textsc{update-generator}(y^{*})$ $\triangleright$ See Section 3.4
  return $\theta^{*},\lambda^{*},v_{r^{*}}^{*},y^{*},z^{*}$
Algorithm 3 Localized Barker-within-Gibbs update for $\pi(v_{r},z,y|v_{s},y_{\nu},\theta,\lambda)$ (see Appendix B).
function update-hidden($\theta,\lambda,v_{s},y,y_{\nu},z$)
  $\nu\sim\kappa(\nu)$ $\triangleright$ S.t. $\nu\subseteq s\mathbin{\backslash}\{0,\omega\}$ and $\operatorname{\mathrm{Pr}}\left[\dot{s}\in\nu\right]<1$ for $\dot{s}\in{s}$. For independence sampler, set $\nu=\emptyset$
[…]
  for $(\dot{s}\sim\ddot{s})\in s$ do
   $v_{r^{\dagger}\cap(\dot{s},\ddot{s})}^{\dagger}\sim\kappa(v_{r^{\dagger}\cap(\dot{s},\ddot{s})}^{\dagger}|v_{\{\dot{s},\ddot{s}\}},y_{[\dot{s},\ddot{s}]}^{\dagger})$ $\triangleright$ See Algorithm 1
  for $(\dot{\tau}\sim\ddot{\tau})\in\tau^{\dagger}$ do
   $z_{(\dot{\tau},\ddot{\tau})}^{\dagger}\sim\mathbb{B}_{(\dot{\tau},\ddot{\tau})}$
   $c_{1}^{(\dot{\tau},\ddot{\tau})}\leftarrow d_{\theta}(z_{(\dot{\tau},\ddot{\tau})}^{\dagger},y_{\dot{\tau}}^{\dagger},v_{\{\dot{\tau},\ddot{\tau}\}}^{\dagger})$
   $\varpi_{1}^{(\dot{\tau},\ddot{\tau})}\leftarrow\textsc{construct-hidden-coin}(\theta,y_{\dot{\tau}}^{\dagger},v_{\{\dot{\tau},\ddot{\tau}\}}^{\dagger})$ $\triangleright$ See Section 3.1
  for $(\dot{\tau}\sim\ddot{\tau})\in\tau$ do
   $c_{2}^{(\dot{\tau},\ddot{\tau})}\leftarrow d_{\theta}(z_{(\dot{\tau},\ddot{\tau})},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$
   $\varpi_{2}^{(\dot{\tau},\ddot{\tau})}\leftarrow\textsc{construct-hidden-coin}(\theta,y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$ $\triangleright$ See Section 3.1
  for $(\dot{\nu}\sim\ddot{\nu})\in s\cap(\{0,\omega\}\cup\{t:y_{t}=y_{t}^{\dagger}\})$ do
   $c_{1}\leftarrow\prod_{(\dot{s}\sim\ddot{s})\in s\cap[\dot{\nu},\ddot{\nu}]}\kappa(v_{r\cap(\dot{s},\ddot{s})}|v_{\{\dot{s},\ddot{s}\}},y_{[\dot{s},\ddot{s}]})\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau^{\dagger}\cap[\dot{\nu},\ddot{\nu}]}c_{1}^{(\dot{\tau},\ddot{\tau})}$
   $c_{2}\leftarrow\prod_{(\dot{s}\sim\ddot{s})\in s\cap[\dot{\nu},\ddot{\nu}]}\kappa(v_{r^{\dagger}\cap(\dot{s},\ddot{s})}^{\dagger}|v_{\{\dot{s},\ddot{s}\}},y_{[\dot{s},\ddot{s}]}^{\dagger})\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau\cap[\dot{\nu},\ddot{\nu}]}c_{2}^{(\dot{\tau},\ddot{\tau})}$
   $\varpi_{1}\leftarrow\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau^{\dagger}\cap[\dot{\nu},\ddot{\nu}]}\varpi_{1}^{(\dot{\tau},\ddot{\tau})}$
   $\varpi_{2}\leftarrow\prod_{(\dot{\tau}\sim\ddot{\tau})\in\tau\cap[\dot{\nu},\ddot{\nu}]}\varpi_{2}^{(\dot{\tau},\ddot{\tau})}$
   $\alpha_{(\dot{\nu},\ddot{\nu})},z_{(\dot{\nu},\ddot{\nu})}^{*}\leftarrow\textsc{2-coin-sep}(c_{1},c_{2},\varpi_{1},\varpi_{2},z_{(\dot{\nu},\ddot{\nu})}^{\dagger},z_{(\dot{\nu},\ddot{\nu})})$
   if $\alpha_{(\dot{\nu},\ddot{\nu})}$ then
     $v_{r^{*}\cap(\dot{\nu},\ddot{\nu})}^{*},y_{[\dot{\nu},\ddot{\nu}]}^{*}\leftarrow v_{r^{\dagger}\cap(\dot{\nu},\ddot{\nu})}^{\dagger},y_{[\dot{\nu},\ddot{\nu}]}^{\dagger}$
   else
     $v_{r^{*}\cap(\dot{\nu},\ddot{\nu})}^{*},y_{[\dot{\nu},\ddot{\nu}]}^{*}\leftarrow v_{r\cap(\dot{\nu},\ddot{\nu})},y_{[\dot{\nu},\ddot{\nu}]}$
  $\textsc{adapt-hidden-proposal}(\{\alpha_{(\dot{\nu},\ddot{\nu})}:(\dot{\nu}\sim\ddot{\nu})\in\nu\})$ $\triangleright$ See Section B
  return $v_{r^{*}}^{*},z^{*},y^{*}$
Algorithm 4 Random walk Barker-within-Gibbs update for $\pi(\theta|v_{\tau},z,y)$ (see Section 3.3).
function update-param($\theta,v_{\tau},z,y$)
  $\theta^{\dagger}\sim\kappa(\theta^{\dagger}|\theta)$ $\triangleright$ Typically (tuned) random walk proposal
  for $(\dot{\tau}\sim\ddot{\tau})\in\tau$ do
   $c_{1}^{(\dot{\tau},\ddot{\tau})}\leftarrow\{\kappa(\theta|\theta^{\dagger})\pi(\theta^{\dagger})\}^{1/(|\tau|-1)}h_{\theta^{\dagger}}(y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$
   $c_{2}^{(\dot{\tau},\ddot{\tau})}\leftarrow\{\kappa(\theta^{\dagger}|\theta)\pi(\theta)\}^{1/(|\tau|-1)}h_{\theta}(y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$
   $\varpi_{1}^{(\dot{\tau},\ddot{\tau})}\leftarrow\textsc{construct-param-coin}(\theta^{\dagger},\theta,y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$ $\triangleright$ See Section 3.1
   $\varpi_{2}^{(\dot{\tau},\ddot{\tau})}\leftarrow\textsc{construct-param-coin}(\theta,\theta^{\dagger},y_{\dot{\tau}},v_{\{\dot{\tau},\ddot{\tau}\}})$ $\triangleright$ See Section 3.1
  $\alpha,z^{*}\leftarrow\textsc{dcbf}(\{c_{1}^{(\dot{\tau},\ddot{\tau})},c_{2}^{(\dot{\tau},\ddot{\tau})},\varpi_{1}^{(\dot{\tau},\ddot{\tau})},\varpi_{2}^{(\dot{\tau},\ddot{\tau})},z_{(\dot{\tau},\ddot{\tau})}:(\dot{\tau}\sim\ddot{\tau})\in\tau\})$
  if $\alpha$ then
   $\theta^{*}\leftarrow\theta^{\dagger}$
  else
   $\theta^{*}\leftarrow\theta$
  $\textsc{adapt-param-proposal}(\alpha,\theta,\theta^{\dagger})$ $\triangleright$ See e.g. [11]
  return $\theta^{*},z^{*}$
Algorithm 5 2-coin algorithm with separate retrospective sampling (see Section 3.1).
function 2-coin-sep($c_{1},c_{2},\varpi_{1},\varpi_{2},z_{1},z_{2}$)
  $z_{1}^{*}\leftarrow z_{1}$
  $z_{2}^{*}\leftarrow z_{2}$
  loop
   $f_{0}\sim\operatorname{\mathrm{Bern}}\left[c_{1}/(c_{1}+c_{2})\right]$
   if $f_{0}$ then
     $f_{1},z_{1}^{*}\leftarrow\varpi_{1}(z_{1}^{*})$
     if $f_{1}$ then return $1,z_{1}^{*}$
   else
     $f_{2},z_{2}^{*}\leftarrow\varpi_{2}(z_{2}^{*})$
     if $f_{2}$ then return $0,z_{2}^{*}$
Algorithm 6 Divide-and-conquer Bernoulli factory (See [44]).
[…]
  if depth = 0 then
[…]
  $\{z_{i}^{*}\}_{i=1}^{n}\leftarrow\{z_{i}\}_{i=1}^{n}$
  loop
   $\alpha_{1},\{z_{i}^{*}\}_{i=1}^{\lfloor n/2\rfloor}\leftarrow\textsc{dcbf}(\{c_{1}^{(i)},c_{2}^{(i)},\varpi_{1}^{(i)},\varpi_{2}^{(i)},z_{i}^{*}\}_{i=1}^{\lfloor n/2\rfloor},\textsc{depth}-1)$
   $\alpha_{2},\{z_{i}^{*}\}_{i=\lceil n/2\rceil}^{n}\leftarrow\textsc{dcbf}(\{c_{1}^{(i)},c_{2}^{(i)},\varpi_{1}^{(i)},\varpi_{2}^{(i)},z_{i}^{*}\}_{i=\lceil n/2\rceil}^{n},\textsc{depth}-1)$
   if $\alpha_{1}=\alpha_{2}$ then
     return $\alpha_{1}$
Algorithm 7 2-coin algorithm with common retrospective sampling (see Section 3.1).
function 2-coin-com($c_{1},c_{2},\varpi_{1},\varpi_{2}$)
  loop
   $f_{0}\sim\operatorname{\mathrm{Bern}}\left[c_{1}/(c_{1}+c_{2})\right]$
   if $f_{0}$ then
     $f_{1},z^{*}\leftarrow\varpi_{1}(z^{*})$
     if $f_{1}$ then return $1,z^{*}$
   else
     $f_{2},z^{*}\leftarrow\varpi_{2}(z^{*})$
     if $f_{2}$ then return $0,z^{*}$